Q1. Solve the inequalities and show the solution sets on the real line.

a) βˆ’2π‘₯2βˆ’π‘₯+3β‰₯0

b) 13π‘₯+23<56π‘₯βˆ’1

c) 3π‘₯βˆ’3β‰₯π‘₯βˆ’5

d) 4<π‘₯2<9

e) √(π‘₯βˆ’1)2 <√4


1
Expert's answer
2021-04-13T12:51:42-0400

a)

-2xΒ²-x+3β‰₯0

=> 2xΒ²+x-3≀0

=> 2xΒ²+3x-2x-3≀0

=> (2x+3)(x-1)≀0

=> (x+1.5)(x-1)≀0

So x "\isin" [-1.5, 1]



b)

13x+23 < 56x-1

=> 13x-56x < -1-23

=> -43x < -24

=> x > 24/43

So x "\isin ( \frac{24}{43} , ∞)"



c)

3x-3β‰₯x-5

=> 3x - x β‰₯ -5+3

=> 2x β‰₯ -2

=> x β‰₯ -1

So x "\isin [-1, ∞)"



d)

4 < xΒ² < 9

=> xΒ² > 4 and xΒ² <9

=> {x <-2 or x > 2 } and -3<x<3

=> {-3<x<-2} or {2<x<3}

So x "\isin (-3,-2)\bigcup(2,3)"



e)

"\sqrt{(x-1)²} < \sqrt{4}"

=> | x - 1 | < 2

=> -2 < x - 1 < 2

=> -2 +1 < x < 2+1

=> -1 < x < 3

So x "\isin (-1, 3)"